OpenStudy (anonymous):

Jackson saves $2 during Month 1. During each subsequent month, he plans to save 3 times the amount saved the previous month. After 9 months, how much money will he have saved in all? A. $16,482 B. $18,386 C. $19,682 D. $20,464

OpenStudy (lieutenantgeneral):

D. 20,464

OpenStudy (anonymous):

how did u get that @LieutenantGeneral , please write the formula for the series

OpenStudy (lieutenantgeneral):

i didn't my bro did

OpenStudy (anonymous):

I get $13122 which is none of the above answers ???

OpenStudy (anonymous):

my formula is 2 * 3 ^ (m-1) where m = month so for month = 1 2 3 .... 9 saving = 2 6 18 .... 13122

OpenStudy (anonymous):

Same here which threw me off

OpenStudy (texaschic101):

2 -- month 1 2(3) = 6 -- month 2 6(3) = 18 -- month 3 18(3) = 54 -- month 4 54(3) = 162 -- month 5 162(3) = 486 -- month 6 486(3) = 1458 -- month 7 1458(3) = 4374 -- month 8 4374(3) = 13122 -- month 9 in total : 13122 + 4374 + 1458 + 486 + 162 + 54 + 18 + 6 + 2 = 19682 (C) I am sure there is an easier way to do this, I just don't know it

OpenStudy (texaschic101):

you could probably use a geometric series formula to solve this

OpenStudy (anonymous):

@texaschic101 u too good!

OpenStudy (texaschic101):

not really....there is an easier way...just not sure. What if the question asked for the 100th month ? I would be in trouble then...lol